OpenStudy (anonymous):

How do you solve 5x-35y? Using distributive properties

Parth (parthkohli):

I guess you want to factorize.

OpenStudy (anonymous):

How do I do it

OpenStudy (campbell_st):

well you can't solve it as the expression 5x - 35y isn't equal to anything. so perhaps you mean factorise...

OpenStudy (campbell_st):

look for a common factor in 5x and 35y its a number

OpenStudy (anonymous):

It says: use distributive property to rewrite the expressi. 5x-35y

OpenStudy (anonymous):

Expression

OpenStudy (campbell_st):

ok... so can you see a number that is a factor of 5x and 35y...?

OpenStudy (anonymous):

5

OpenStudy (campbell_st):

yep great so if you divide 5x by 5 what do you get same with 35 y divided by 5 what do you get?

OpenStudy (anonymous):

X 7y

Parth (parthkohli):

Very good! Separate with minus.

OpenStudy (campbell_st):

yep so the common factor is outside brackets with the x and 7y inside the operator is unchanged so what do you think it would look like..?

OpenStudy (anonymous):

So X-7y

OpenStudy (campbell_st):

that bit is inside brackets... and 5 is outside... what does it look like..?

OpenStudy (anonymous):

There are no brackets in the problem

OpenStudy (anonymous):

X (-7y)

Parth (parthkohli):

When you factor, you get brackets in the end even when the problem doesn't involve them. For example, \(2x + 2y = 2(x + y)\)

OpenStudy (campbell_st):

not quite common factor of 5 x - 7y inside brackets 5x - 35y = 5(x - 7y)

OpenStudy (anonymous):

Ok thx ppl